A small 10^{9}kg asteroid, moving in the positive x direction at 10,000 m/s, collides with a 10^{8}kg asteroid moving in the positive y direction at 20,000 m/s. The two stick together and form a single object.

What is the velocity, and angle of the pair after the collision?
Answer #1

using momentum conservation:

In the x axis:

m_1v_1=(m_1+m_2)vcos\theta (I)

In the y axis:

m_2v_2=(m_1+m_2)vsin\theta  (II)

(II)/(I)

tan\theta=\frac{m_2v_2}{m_1v_1}=\frac{10^{8}\cdot 20000}{10^{9}\cdot 10000}

\theta=11.30^{\circ}

Plugging the angle in any of the equations:

m_1v_1=(m_1+m_2)vcos\theta

10^{9}\cdot 10000=(10^{9}+10^{8})vcos11.3

solving for v

v=9270.62\, m/s